# MACE-OFF23: Transferable Machine Learning Force Fields for Organic Molecules

### v2

@@ Abstract
Classical empirical force fields have dominated biomolecular simulation for over 50 years. Although widely used in drug discovery, crystal structure prediction, and biomolecular dynamics, they generally lack the accuracy and transferability required for predictive modelling. In this paper, we introduce MACE-OFF23, a transferable force field for organic molecules created using state-of-the-art machine learning technology and first-principles reference data computed with a high level of quantum mechanical theory. MACE-OFF23 demonstrates the remarkable capabilities of local, short-range models by accurately predicting a wide variety of gas and condensed phase properties of molecular systems. It produces accurate, easy-to-converge dihedral torsion scans of unseen molecules, as well as reliable descriptions of molecular crystals and liquids, including quantum nuclear effects. We further demonstrate the capabilities of MACE-OFF23 by determining free energy surfaces in explicit solvent, as well as the folding dynamics of peptides. Finally, we simulate a fully solvated small protein, observing accurate secondary structure and vibrational spectrum. These developments enable first-principles simulations of molecular systems for the broader chemistry community at high accuracy and low computational cost.
@@ I Introduction
In contrast, simulating bio-organic systems entails a different set of trade-offs, with greater emphasis on simulating systems over long timescales. This means that empirical force fields, which sacrifice accuracy for computational speed, continue to be used routinely to study molecular liquids, crystals, biological systems, and drug-like molecules [19, 20, 21, 22].
@@
Two alternatives to empirical force fields are available. The first is semi-empirical quantum mechanics, such as the series of extended tight-binding models [23], which represents a low-cost solution for small molecules. The method is limited by its moderate accuracy compared to quantum chemistry methods, its restriction to modelling non-periodic systems, and its cubic scaling with system size.
@@
Secondly, a number of transferable machine learning force fields have also been developed for organic chemistry. The most notable are the series of ANI [24, 25, 26, 27] and AIMNet potentials [28, 29, 8]. ANI potentials pioneered the use of local symmetry function-based feedforward neural networks [30] trained on a large dataset of organic molecular geometries [31, 32] to create transferable force fields. The ANI-2x model became the most widely adopted ML force field and therefore serves as one of the primary points of comparison in this paper. The ANI-2x model was recently combined with a polarizable electrostatic model [33] in a hybrid ML/MM simulation setting, and also with a neural network based dispersion correction [34]. The AIMNet models apply a message passing architecture [35], where the initial embeddings are the ANI symmetry functions. They have also extended the applicability of the models to a larger set of chemical elements, as well as to charged species. These models relax the locality assumptions by incorporating electrostatic and dispersion interactions. The PhysNet model uses a message passing architecture [35] and in addition to the semi-local terms also includes long-range electrostatic and dispersion interactions [36].
@@
Further recent bio-organic force fields include the FENNIX model, which combines a local equivariant machine learning model with a physical long-range functional form for electrostatics and dispersion [37]. The model was trained to reproduce the CCSD(T)/CBS energies of small molecules and molecular dimers. It was shown that FENNIX can be used to run stable dynamics of liquid water, solvated alanine dipeptides, and an entire protein in the gas phase. However, wider benchmarking of this model is required to assess the accuracy of the intramolecular potential outside the training set and of condensed phase molecular dynamics simulations.
@@
Similarly, the ANA2B potential employs a short-ranged ML potential, with long-ranged, classical multipolar electrostatics, polarization and dispersion interactions [38]. Although this long-ranged model shows promising accuracy for condensed phase properties and crystal structure ranking, its accuracy and computational performance has not yet been demonstrated for larger biomolecules. Finally, the GEMS model [39], built on the SpookyNet architecture [40], is another recent ML force field for biomolecular simulations. Although SpookyNet has demonstrated application to more challenging condensed phase simulations, including protein dynamics, these required a significant number of reference quantum chemistry calculations on relevant large peptide fragments for each new simulation to obtain a stable model and is therefore not transferable in the sense that the previously discussed force fields are.
@@
In this paper, we introduce a series of three purely local transferable bio-organic machine learning force fields called MACE-OFF23. The force fields are parameterized for the ten most important chemical elements for organic chemistry H, C, N, O, F, P, S, Cl, Br, and I. They are capable of accurately describing intra- and intermolecular interactions of neutral closed-shell systems. This enables the simulation of a wide range of chemical systems, from molecular liquids and crystals to drug-like molecules and biopolymers.
@@
The models were validated on a number of tasks, including the prediction of small molecule torsion barriers, geometry optimizations, calculations of lattice parameters and enthalpies of formation of molecular crystals, and the calculation of the Raman spectra of molecular crystals, including nuclear quantum effects. We have also validated the models for predicting the densities and heats of vaporization of a range of molecular liquids. In particular, we study how well MACE-OFF23 reproduces the basic properties of water, including density and radial distribution functions. To further showcase the capabilities of the model, we computed the free energy surface of the alanine tripeptide (Ala3) in vacuum and explicit water, and simulated the folding of Ala15 at different temperatures. We also carried out an all-atom simulation of the protein crambin in explicit water (18 000 atoms), where we observed the expected secondary structure and computed its vibrational spectrum finding good agreement with previously reported experimental data. Finally, we tested the computational speed of the current implementation in the LAMMPS and OpenMM simulation packages.
@@ II.1 The MACE architecture
The MACE model [2] is a force field that maps the positions and chemical elements of atoms to the system’s potential energy. Linear scaling with system size is achieved by decomposing the total energy into atomic site energies. First, a graph is defined by connecting two nodes (atoms) by an edge if they are in each other’s local environment. The local environment $\mathcal{N}(i)$ is the set of all atoms $j$ around the central atom $i$ for which $\|\bm{r}_{ij}\|\leq r_{\text{cut}}$, where $\bm{r}_{ij}$ is the vector from atom $i$ to atom $j$ and $r_{\text{cut}}$ is the cutoff hyperparameter. The array of features of node $i$ is denoted by ${\bm{h}}_{i}^{(t)}$ and is expressed in the spherical harmonic basis, with its elements indexed by $l$ and $m$. The superscript in $t$ indicates the iteration steps (corresponding to “layers” of message passing in the language of graph neural networks). All MACE-OFF23 models are made up of two layers. The ${\bm{h}}^{(t)}_{i}$ features depend on the chemical environment of the atoms.
@@
×$\qquad\qquad\times Y^{m_{1}}_{l_{1}}(\bm{\hat{r}}_{ij})h^{(t)}_{j,kl_{2}m_{2}}$ (2)
where $Y_{l}^{m}$ are the spherical harmonics, and $C_{\eta_{1},l_{1}m_{1}l_{2}m_{2}}^{l_{3}m_{3}}$ denotes the Clebsch-Gordan coefficients. There are multiple ways of constructing an equivariant combination with a given symmetry $(l_{3},m_{3})$, and these multiplicities are enumerated by the path index $\eta_{1}$ [47, 2].
The one-particle basis $\phi$ is summed over the neighborhood and expanded in a linear combination of $k$ channels with learnable weights to form the permutation invariant atomic basis $A_{i}$:
@@
Higher-order (many-body) symmetric features are created on each atom by taking products of the atomic basis, $A$, with itself $\nu$ times, resulting in the “product basis”. The product basis is then contracted with the generalised Clebsch-Gordan coefficients $\mathcal{C}^{LM}_{\eta_{\nu}\bm{l}\bm{m}}$ to obtain the equivariant higher-order basis, $\bm{B}_{i}$ [47, 2]. The maximum body-order is controlled by the parameter $\nu$ and is fixed at 3 (corresponding to 4-body terms, which includes the central atom) for all MACE-OFF23 models:
@@
where bold $\bm{l}\bm{m}$ denostes the $\nu$-tuple of $l$ and $m$ values and similarly to Equation (II.1), $\eta_{\nu}$ enumerates the number of possible couplings to create the features with equivariance $LM$.
@@
The recursive update of the node features ($t:0\rightarrow 2$) is obtained by adding the message to the atoms’ features from the previous iteration, with weights that depend on the chemical element ($z_{i}$) that are also responsible for the mixing of the chemical embedding $k$ channels:
@@
Since the initial node features ${\bm{h}}^{(0)}$ are only dependent on the chemical element of the node, the second term of Eq. (6) is not included in the first layer. This makes it possible to set the energy of the isolated atoms (that is, those with no neighbours) exactly [48], which is often desirable [49, 37].
The MACE models in this article have an effective receptive field of $2\times r_{\text{cut}}$ due to the two message-passing layers used in each model.
@@
The forces and stresses on the atoms are calculated by taking analytical derivatives of the total potential energy with respect to the positions of the atoms.
@@ II.2 Training data
[…]
@@
Table 1: Summary of training and test sets The columns PubChem to Solvated Amino Acids correspond to the SPICE dataset [50].
PubChem
DES370K Monomers
DES370K Dimers [58]
Dipeptides
Solvated Amino Acids
Water
QMugs [59]
Tripeptides
Chemical elements
H, C, N, O, F, P, S, Cl, Br, I
H, C, N, O, F, P, S, Cl, Br, I
H, C, N, O, F, P, S, Cl, Br, I
H, C, N, O, S
H, C, N, O, S
H, O
H, C, N, O, F, P, S, Cl, Br, I
H, C, N, O
System size
3-50
3-22
4-34
26-60
79-96
3-150
51-90
30-69
# Train
646821
16861
263065
19773
948
1597
2748
0
# Test
33884
889
13896
1025
52
84
144
898
@@
The SPICE dataset only contains small molecules of up to 50 atoms. To facilitate the learning of intramolecular non-bonded interactions, we augmented SPICE with larger 50–90 atom molecules randomly selected from the QMugs dataset [59]. The geometries were generated by running molecular dynamics simulations using GFN2-xTB [23] similarly to the protocol described in Ref. [60]. The energies and forces were re-evaluated at the level of QM theory used in SPICE. Finally, to obtain a better description of water, the dataset was further augmented with a number of water clusters carved out of molecular dynamics simulations of liquid water [61], with sizes of up to 50 water molecules.
In addition to the 5% of each subset, part of the COMP6 tripeptide geometry dataset [25] was also recomputed at the SPICE level of theory and used as part of the test evaluation, but not for training.
@@
After training the small MACE model (see Section II.3) and observing the training errors, we noticed the presence of outliers in the dataset. This is probably caused by errors in the underlying electronic structure calculations, some of which have been documented on the SPICE GitHub repository.***https://github.com/openmm/spice-dataset To confirm this, we also re-evaluated a selection of the configurations that had the highest error, using the level of DFT used in SPICE, and found that for about a third of the configurations, the recomputed energies and forces agreed well with the MACE prediction and not with the original DFT labels. To purify the dataset, we removed from the training set the configurations that had a maximum force error greater than 2 eV/Å. This meant the removal of just 808 configurations, many of which contained heavy elements, in particular phosphorus and iodine, that might have a more challenging electronic structure.
II.3 Training details
@@
The MACE model has parameters that enable systematic control of model expressivity (accuracy) against computational cost. In this paper, we present three variants of the MACE-OFF23 model, a small, a medium, and a large one, denoted in the text as MACE-OFF23(S), MACE-OFF23(M) and MACE-OFF23(L) respectively. The models get more accurate with size, but also have an increasing computational cost. The small model is well suited for large-scale simulations, the medium model offers a good balance of speed and accuracy, and the large model is best used for small systems or when the highest possible accuracy is desirable.
Table 2: Hyperparameters of the three MACE-OFF23 models
Small Medium Large
Cutoff radius (Å) 4.5 5.0 5.0
Chemical channels 96 128 192
$k$ (Eq. (1))
max L (Eq. (5)) 0 1 2
The hyperparameters of the models are displayed in Table 2. The MACE-OFF23(S) model has 96 channels, $L=0$ invariant messages, and a cutoff of 4.5 Å. The MACE-OFF23(M) model has 128 channels and $L=1$ messages, and finally the MACE-OFF23(L) has 192 channels and $L=2$ equivariant messages. Medium and large models have a cutoff of 5.0 Å in each layer. All models used two layers and a body order of 4 in each layer. The models were trained using the PyTorch [62] implementation of MACE, available on https://github.com/ACEsuit/mace. For more information about the training, please refer to the SI.

### v3

# MACE-OFF: Short Range Transferable Machine Learning Force Fields for Organic Molecules

@@ Abstract
Classical empirical force fields have dominated biomolecular simulation for over 50 years. Although widely used in drug discovery, crystal structure prediction, and biomolecular dynamics, they generally lack the accuracy and transferability required for first-principles predictive modeling. In this paper, we introduce MACE-OFF, a series of short range transferable force fields for organic molecules created using state-of-the-art machine learning technology and first-principles reference data computed with a high level of quantum mechanical theory. MACE-OFF demonstrates the remarkable capabilities of short range models by accurately predicting a wide variety of gas and condensed phase properties of molecular systems. It produces accurate, easy-to-converge dihedral torsion scans of unseen molecules, as well as reliable descriptions of molecular crystals and liquids, including quantum nuclear effects. We further demonstrate the capabilities of MACE-OFF by determining free energy surfaces in explicit solvent, as well as the folding dynamics of peptides. Finally, we simulate a fully solvated small protein, observing accurate secondary structure and vibrational spectrum. These developments enable first-principles simulations of molecular systems for the broader chemistry community at high accuracy and relatively low computational cost.
@@ I Introduction
In contrast, simulating bio-organic systems entails a different set of trade-offs, with greater emphasis on simulating large systems over long timescales. This means that empirical force fields, which sacrifice accuracy for computational speed, continue to be used routinely to study molecular liquids, crystals, biological systems, and drug-like molecules [19, 20, 21, 22].
@@
Two alternatives to empirical force fields are available. The first is semi-empirical quantum mechanics, such as the series of extended tight-binding models [23], which represents a low-cost solution for small molecules. The method is limited by its moderate accuracy compared to quantum chemistry methods, its restriction to modelling non-periodic systems (if not in principle, but in widely used implementations), and its cubic scaling with system size.
@@
More recently, a number of transferable machine learning force fields have also been developed for organic chemistry. By “transferable” , here we mean that these potentials are capable of generalising in system size and both chemical and conformation space sufficiently well to perform stable molecular dynamics and accurate property prediction for a wide range of molecular systems beyond those on which the model was trained. The most notable are the series of ANI [24, 25, 26, 27] and AIMNet potentials [28, 29, 8]. ANI potentials pioneered the use of local symmetry function-based feedforward neural networks [30] trained on a large dataset of organic molecular geometries [31, 32] to create transferable ML force fields. The ANI-2x model became the most widely adopted ML force field and therefore serves as one of the primary points of comparison in this paper. The ANI-2x model was recently combined with a polarizable electrostatic model [33] in a hybrid ML/MM simulation setting, and also with a neural network based dispersion correction [34]. The AIMNet models apply a message passing architecture [35], where the initial embeddings are the ANI symmetry functions. They have also extended the applicability of the models to a larger set of chemical elements, as well as to charged species. These models relax the locality assumptions by incorporating electrostatic and dispersion interactions. The PhysNet model uses a message passing architecture [35] and in addition to the semi-local terms also includes long-range electrostatic and dispersion interactions [36].
@@
Other recent bio-organic force fields include the FENNIX model, which combines a local equivariant machine learning model with a physical long-range functional form for electrostatics and dispersion [37]. The model was trained to reproduce the CCSD(T)/CBS energies of small molecules and molecular dimers. It was shown that FENNIX can be used to run stable dynamics of liquid water, solvated alanine dipeptides, and an entire protein in the gas phase. However, further benchmarking of this model is required to assess the accuracy of the intramolecular potential outside the training set and of condensed phase molecular dynamics simulations.
@@
Similarly, the ANA2B potential employs a short-ranged ML potential, with long-ranged, classical multipolar electrostatics, polarization and dispersion interactions [38]. Although this long-ranged model shows promising accuracy for condensed phase properties and crystal structure ranking, its accuracy and computational performance has not yet been demonstrated for larger biomolecules.
The GEMS model [39], built on the SpookyNet architecture [40], is another recent ML force field for biomolecular simulations. Whilst a transferable version of this potential is capable of producing stable dynamics for challenging biomolecular systems, certain properties, for example the folding dynamics of small peptides, require the addition of system-specific training configurations in order to accurately reproduce experimental properties. More recently, the transferable SO3LR potential was introduced by the same group of researchers [41]. This is based on the equivariant SO3krates architecture and also includes analytic dispersion and electrostatic long range contributions. Whilst this potential is capable of fast and stable simulations of biomolecular systems, we are not yet aware of more extensive quantitative benchmarking on off-equilibrium energetics or thermodynamic properties.
@@
In this paper, we introduce a family of five purely local (i.e. short range), transferable bio-organic machine learning force fields which we refer to collectively as MACE-OFF. The force fields are parameterized for the ten most important chemical elements for organic chemistry: H, C, N, O, F, P, S, Cl, Br, and I. They are capable of accurately describing intra- and intermolecular interactions of neutral closed-shell systems. This enables the simulation of a wide range of chemical systems, from molecular liquids and crystals to drug-like molecules and biopolymers.
@@
The models were validated on a number of tasks, including the prediction of small molecule torsion barriers, geometry optimizations, calculations of lattice parameters and enthalpies of formation of molecular crystals, and the calculation of the Raman spectra of molecular crystals, including nuclear quantum effects. We have also validated the models for predicting the densities and heats of vaporization of a range of molecular liquids. In particular, we study how well MACE-OFF reproduces the fundamental properties of water, including density and radial distribution functions. To further showcase the capabilities of the model, we computed the free energy surface and J-coupling parameters of alanine tripeptide (Ala3) in vacuum and explicit water, and simulated the folding of Ala15 at different temperatures. We also carried out an all-atom simulation of the protein crambin in explicit water (18 000 atoms), where we observed the correct secondary structure motifs and computed its vibrational spectrum, finding good agreement with other ML potentials and previously reported experimental data. Finally, we tested the computational speed of the current implementation in the LAMMPS and OpenMM simulation packages, demonstrating both strong and weak scaling.
While truly general purpose force fields for biomolecular modeling do need to include formally long-range Coulomb interactions, the purpose of this paper is to push the capabilities of short-range potentials as far as they can go, in part as preparation for including explicit polarizable electrostatic interactions in the future.
@@ II.1 The MACE architecture
The MACE model [2] is a force field that maps the positions and chemical elements of atoms to the system’s potential energy. Linear scaling with system size is achieved by decomposing the total energy into atomic site energies. First, a graph is defined by connecting two nodes (atoms) by an edge if they are in each other’s local environment. The local environment $\mathcal{N}(i)$ is the set of all atoms $j$ around the central atom $i$ for which $\|\bm{r}_{ij}\|\leq r_{\text{cut}}$, where $\bm{r}_{ij}$ is the vector from atom $i$ to atom $j$ and $r_{\text{cut}}$ is the cutoff hyperparameter. The array of features of node $i$ is denoted by ${\bm{h}}_{i}^{(t)}$ and is expressed in the spherical harmonic basis, with its elements indexed by $l$ and $m$. The superscript in $t$ indicates the iteration steps (corresponding to “layers” of message passing in the language of graph neural networks). All MACE-OFF models are made up of two layers. The ${\bm{h}}^{(t)}_{i}$ features depend on the chemical environment of the atoms for $t>0$.
@@
$\qquad\qquad\times Y^{m_{1}}_{l_{1}}(\bm{\hat{r}}_{ij})h^{(t)}_{j,kl_{2}m_{2}}$ (2)
where $Y_{l}^{m}$ are the spherical harmonics, and $C_{\eta_{1},l_{1}m_{1}l_{2}m_{2}}^{l_{3}m_{3}}$ denotes the Clebsch-Gordan coefficients. There are multiple ways of constructing an equivariant combination with a given symmetry $(l_{3},m_{3})$, and these multiplicities are enumerated by the index $\eta_{1}$ [48].
The one-particle basis $\phi$ is summed over the neighborhood and the $k$ channels are mixed together with learnable weights to form the permutation invariant atomic basis $A_{i}$:
@@
Higher-order (many-body) symmetric features are created on each atom $i$ by taking tensor products of the atomic basis, $A_{i}$, with itself $\nu$ times, resulting in the “product basis”. The product basis is then contracted with the generalized Clebsch-Gordan coefficients, $\mathcal{C}^{LM}_{\eta_{\nu}\bm{l}\bm{m}}$, to obtain the equivariant many-body basis, $\bm{B}_{i}$ [48],
@@
where bold $\bm{l}\bm{m}$ denotes the $\nu$-tuple of $l$ and $m$ values and similarly to Equation (II.1), $\eta_{\nu}$ enumerates the number of possible couplings to create the features with equivariance $LM$. The maximum body-order is controlled by the parameter $\nu$ and is fixed at 3 (corresponding to 4-body terms, which includes the central atom) for all MACE-OFF models in this work.
@@
The recursive update of the node features ($t:0\rightarrow 2$) is obtained by adding the message to the atoms’ features from the previous iteration, with weights that depend on the chemical element ($z_{i}$):
@@
Since the initial node features ${\bm{h}}^{(0)}$ are only dependent on the chemical element of the atom, the second term of Eq. (6) is not included in the first layer. This makes it possible to set the energy of the isolated atoms (that is, those with no neighbours) exactly [49], which is often desirable [50, 37]. The MACE models in this work have an effective receptive field of $2\times r_{\text{cut}}$ due to the two layers of message passing.
@@
The forces and stresses on the atoms are calculated by taking analytical derivatives of the total potential energy with respect to the positions of the atoms and the components of the deformation tensor in the periodic setting, as is usual.
@@ II.2 Training data
The core of the training set is version 1 of the SPICE dataset [51] with 95% of the data used for training and validation, and 5% set aside for testing. Test/train splitting was performed at the molecule level, ensuring conformers of the same molecule do not appear in both train/validation and test sets. Table 1 summarizes the training and test sets. The MACE-OFF models are trained to reproduce the energies and forces computed at the ωB97M-D3(BJ)/def2-TZVPPD level of quantum mechanics [52, 53, 54, 55, 56], as implemented in the PSI4 software package [57]. We have used a subset of SPICE that contains the ten chemical elements H, C, N, O, F, P, S, Cl, Br, and I, and has a neutral formal charge (see Table 2 in Ref. [51] for element coverage in the dataset). We have also removed the ion pairs subset. Overall, we used about 85% of the full SPICE dataset. The geometries in the SPICE dataset have been generated by running molecular dynamics simulations using classical force fields [58], performed at both 300K and 500K, and sampling maximally different conformations from the resulting trajectories [51].
@@
Table 1: Summary of training and test sets. The columns PubChem to Solvated Amino Acids correspond to the SPICE dataset [51].
PubChem DES370K Monomers DES370K Dimers [59] Dipeptides Solvated Amino Acids Water QMugs [60] Tripeptides
Chemical elements H, C, N, O, F, P, S, Cl, Br, I H, C, N, O, F, P, S, Cl, Br, I H, C, N, O, F, P, S, Cl, Br, I H, C, N, O, S H, C, N, O, S H, O H, C, N, O, F, P, S, Cl, Br, I H, C, N, O
System size 3-50 3-22 4-34 26-60 79-96 3-150 51-90 30-69
# Train 646821 16861 263065 19773 948 1597 2748 0
# Test 33884 889 13896 1025 52 84 144 898
@@
The SPICE dataset only contains small molecules of up to 50 atoms. To facilitate the learning of intramolecular non-bonded interactions, we augmented SPICE with larger 50–90 atom molecules randomly selected from the QMugs dataset [60]. The geometries were generated by running molecular dynamics simulations using GFN2-xTB [23] similarly to the protocol described in Ref. [61]. The energies and forces were re-evaluated at the level of QM theory used in SPICE. Finally, to obtain a better description of water, the dataset was further augmented with a number of water clusters carved out of molecular dynamics simulations of liquid water [62], with sizes of up to 50 water molecules. In addition to the 5% of each subset, part of the COMP6 tripeptide dataset [25] was also recomputed at the SPICE level of theory and used as part of the test evaluation, but not for training.
@@
After fitting the small MACE model (see Section II.3) and observing the training errors, we noticed the presence of outliers in the dataset. This is probably caused by errors in the underlying electronic structure calculations, some of which have been documented on the SPICE GitHub repository.***https://github.com/openmm/spice-dataset To confirm this, we also re-evaluated a selection of the configurations that had the highest error, using the level of DFT used in SPICE, and found that for about a third of the configurations, the recomputed energies and forces agreed well with the MACE prediction and not with the original DFT labels. To purify the dataset, we removed from the training set the configurations that had a maximum force error greater than 2 eV/Å. This meant the removal of just 808 configurations, many of which contained heavy elements, in particular phosphorus and iodine, that might have a more challenging electronic structure.
In order to understand the role additional data plays in the performance of the model, we also investigated the difference in performance due to inclusion of additional configurations released as part of the second version of SPICE [63]. The updated dataset, which we incorporate into the MACE-OFF24(M) model (see Section II.3), contains among others, an additional 208,000 configurations comprising solvated PubChem molecules and amino acid-ligand pairs. The former were sampled from molecular dynamics using the same protocol as in the original dataset, whilst the amino acid-ligand pairs were extracted from structures in the PDB. This release also contained 1000 water clusters, containing up to 90 atoms. These were also included and used to complement the set of water clusters we had already included in the extended dataset.
In common with the first version of SPICE, close inspection of this new set revealed several nonphysical geometries with correspondingly high DFT forces. These configurations were removed by applying a maximum force filter of 15 eV/Å. We also observed that many phosphorous-containing compounds were present among the outliers, as well as configurations with highly distorted geometries, which also suggest challenging electronic structure. We also found that many configurations had forces that do not sum to zero. To avoid the possibility of these configurations poisoning the dataset, we applied a total force filter of 0.1 eV/Å. In total this removed 10,372 configurations from the SPICE v2 subsets.
II.3 Model details
@@
The MACE model has parameters that enable systematic control of model expressivity (and subsequent accuracy) against computational cost. In this paper, we present three variants of the MACE-OFF23 model, a small, a medium, and a large one, denoted in the text as MACE-OFF23(S), MACE-OFF23(M) and MACE-OFF23(L) respectively. The models get more accurate with size, but also have an increasing computational cost. The small model is well suited for large-scale simulations, the medium model offers a good balance of speed and accuracy, and the large model is best used for small systems or when the highest possible accuracy is desirable.
Table 2: Hyperparameters of the MACE-OFF models
23(S) 23(M) 23(M)b 23(L) 24(M)
Cutoff radius (Å) 4.5 5.0 6.0 5.0 6.0
Chemical channels 96 128 128 192 128
$k$ (see Eq. (1))
max L (see Eq. (5)) 0 1 1 2 1
SPICE version 1 1 1 1 2
The hyperparameters of the models are displayed in Table 2. The MACE-OFF23(S) model has 96 channels, $L=0$ invariant messages, and a cutoff of 4.5 Å. The MACE-OFF23(M) model has 128 channels and $L=1$ messages, and the MACE-OFF23(L) has 192 channels and $L=2$ equivariant messages. Medium and large models have a cutoff of 5.0 Å in each layer. In light of the additional configurations added to the SPICE dataset during this work, we present two additional models that improve upon the performance of the original three MACE-OFF23 models. MACE-OFF23(M)b is identical to the medium model except for the increased 6 Å cutoff, which we find is crucial for recovering certain properties of condensed phase systems. MACE-OFF24(M) retains the extended 6 Å cutoff and includes additional configurations from SPICE v2 dataset [63]. MACE-OFF23(M)b can therefore be viewed as an ablation of the MACE-OFF24(M) model, enabling insight into which properties are driven by the extended cutoff and those that additionally require the additional nonbonded interaction datasets. All models were trained using the PyTorch [64] implementation of MACE, available on https://github.com/ACEsuit/mace. For more information about the training, please refer to the SI.
II.4 MACE-OFF23-$\mu$
In addition to the series of MACE-OFF force field models, we also present MACE-OFF23-$\mu$, which predicts the total dipole moment of molecules and can be used to estimate vibrational spectra based on the autocorrelation of the dipole moment. The model architecture is identical to the MACE architecture described above, with the only difference being the readout function, which produces an $L=1$ tensor for each atom instead of the atomic site energies. Given the equivariant order of the readout, the smallest possible model corresponds to the ‘medium’ architecture, which is the one considered in this work. Other than the readout’s equivariant order, all hyperparameters match those of the MACE-OFF23(M) model. MACE-OFF23-$\mu$ was trained on the subset of SPICE v1 for which a total dipole moment was available, resulting in the exclusion of the QMugs and water subsets.